HUBBLE SPACE TELESCOPE DAILY REPORT #1047 PERIOD COVERED: 0700 EST Daily Status Report as of 006/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed S/C FOC/96 4735 (COSTAR/FOC Phase Retrieval) The COSTAR and the Faint Object Camera were used to continue to obtain data required for phase retrieval. A long series of out of focus images was taken of standard target GRW70D5824. The first scheduled corrective SIP maneuver was not needed. After the first set of observations was complete, the second planned SIP corrective maneuver was successfully uplinked to the spacecraft. These observations are intended to find the best position for the FOC focus and mirror tilt. The observations were reported to look satisfactory. 1.2 Completed WF/PC-2 4721 (WF/PC-2 Decontamination and TEC Turn On) The WF/PC-2 was used to make ultraviolet photometric observations of the star standard A+81D266, plus K-Spots, and dark calibration images. All four sets of CCDs were used. A total of 18 images were received and displayed in the OSS. Two images were reported to be saturated in the core, and the other images were reported to be nominal. 1.3 Completed COSTAR/FOC/96 4734 (COSTAR/FOC Fine Alignment) The Faint Object Camera (f/96) was used for a Mode 1 target acquisition to center the target star HZ21. A 3X3 COSTAR tilt scan centered on the best FOC position was completed. A 256X256 FOC image was taken at each of nine points, followed by a COSTAR fine focus sweep. These operations should find the best position of focus and tilt. The observations were monitored in the OSS, and the data quality was assessed. No anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 4 Successful : 4 Scheduled Reacquisitions : 4 Successful : 4 2.2 FHST Updates: Scheduled :4 Successful :3 FHST#1 and 2 failed to perform a full maneuver update at 005/1201Z. FHST #2 failed three error box iterations. 2.3 Significant HSTARs: None 2.4 Operations Notes: SMOV activities continue on or slightly ahead of the planned timeline. All observatory subsystems are performing nominally. 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete FGS 5063 (Linking the HIPPARCOS to the Extragalactic Reference Frame, Part 5 of 6) 3.2 Complete WF/PC-2 5193 (Observations of an External Source) 3.3 Complete WF/PC-2 5195 (Observations of an External Source) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: None /DMH -END-